# Complex exponential integral means spectra of univalent functions and the Brennan conjecture

### v8

@@ Abstract.
In this paper we investigate the complex exponential integral means spectra of univalent functions in the unit disk. We show that all integral means spectrum (IMS) functionals for complex exponents on the universal Teichmüller space, the closure of the universal Teichmüller curve, and the universal asymptotic Teichmüller space are continuous. We also show that the complex exponential integral means spectrum of any univalent function admitting a quasiconformal extension to the extended complex plane is strictly less than the universal integral means spectrum. These extend some related results in our recent work [38]. Here we employ a different and more direct approach to prove the continuity of IMS functional on the universal asymptotic Teichmüller space. Additionally, we completely determine the integral means spectra of all univalent rational functions in the unit disk. As a consequence, we show that the Brennan conjecture is true for this class of univalent functions. Finally, we present some remarks and raise some problems and conjectures regarding IMS functionals on Teichmüller spaces, univalent rational functions, and a multiplication operator whose norm is closely related to the Brennan conjecture.
@@
Complex exponential integral means spectrum of univalent function, universal integral mean spectrum, Brennan conjecture, univalent function with a quasiconformal extension, univalent rational function, univalent polynomial, universal Teichmüller space, universal asymptotic Teichmüller space, universal Teichmüller curve
@@ 1. Introduction
For real $\tau$, the integral means measure boundary expansion and compression associated with a given conformal map. Allowing $\tau$ to be a complex exponent, we should also take into account the rotation.
The universal integral mean spectra $B(\tau)$ and $B_{b}(\tau)$ are defined as
@@
Because of the important relation to the fine properties of harmonic measures in the complex plane, see [43], the universal integral means spectra of univalent functions have been studied extensively in recent years. It is an important and difficult problem to find out the exact values of the universal integral mean spectra $B(\tau)$ and $B_{b}(\tau)$. We shall first review some main known results and open problems for real $\tau$ in this topic. It was first observed by Makarov in [43] that
@@
(1) For fixed $\tau\in\mathbb{C}$ with $\tau\neq 0$, what topology can be endowed upon the class $\mathcal{S}_{q}$ such that the integral means spectrum $\beta_{f}(\tau)$ is continuous on $\mathcal{S}_{q}$? (2) For each $\tau\neq 0$, does there exist at least one extremal function for $B_{b}(\tau)$? (3) If the extremal functions for $B_{b}(\tau)$ exist, in which subset of $\mathcal{S}$ do they lie?
Here and later, we say a function $f\in\mathcal{S}$ is an extremal function for $B_{b}(\tau)$ if $\beta_{f}(\tau)=B_{b}(\tau)$. In the recent work [38], we have studied Problem 1.11 for real $\tau$. It has been pointed out that the functional $I:f\rightarrow\beta_{f}(\tau),f\in\mathcal{S}_{q}$ is not continuous on $\mathcal{S}_{q}$ under the locally uniformly convergence topology, see [38, Remark 5.8]. In [38], we introduced and studied the integral means spectrum (IMS) functionals on the Teichmüller spaces. It has been proved in [38] that all IMS functionals on the universal Teichmüller space, the closure of the universal Teichmüller curve, and the universal asymptotic Teichmüller space are continuous. It was also shown in [38] that the real exponential integral means spectrum of any univalent function admitting a quasiconformal extension to $\widehat{\mathbb{C}}$ is strictly less than the universal integral means spectrum. This result means that, for $\tau\in\mathbb{R}$ with $\tau\neq 0$, the extremal functions for $B_{b}(\tau)$ cannot be from the class $\mathcal{S}_{q}$. The first main purpose of this paper is to extend these results to the complex setting. We shall prove that
@@ 4. Proof of Theorem 1 and 2
Note that $d_{AT}([\mu]_{AT},[\nu]_{AT})\leq d_{T}([\mu]_{T},[\nu]_{T})$ for any $\mu,\nu\in M(\Delta^{*})$, we see that the statement that $I_{AT}$ is continuous on $AT$ implies Theorem 1. Hence we only need to prove Theorem 2. By studying the Pre-Schwarzian derivative model of the universal asymptotic Teichmüller space, we have proved in [38] that Theorem 2 holds for real $\tau$. We will use a more direct way to show that Theorem 2 holds for all complex number $\tau$. More precisely, we will prove Theorem 2 by using some arguments from the theory of Teichmüller spaces and some results from the extremal theory of quasiconformal mappings.
@@ 4.1. Isomorphism mapping between Teichmüller spaces
Here $\chi=\frac{\overline{\partial f_{\omega}}}{\partial f_{\omega}}$. We next will write $\mu^{\sharp}(\zeta)=\sharp(\mu)(\zeta)$. Then, from $f_{\mu}=f_{\mu^{\sharp}}\circ f_{\omega}$, we obtain
@@ 4.2. Extremal theory of quasiconformal mappings
It is known that $H_{0}([\mu]_{AT(\Omega)})=h_{0}([\mu]_{T(\Omega)})$. We say some $\widetilde{\mu}\in[\mu]_{AT({\Omega})}$ is non-Strebel extremal in $\mu]_{AT({\Omega})}$ if it satisfies that $H_{0}([\mu]_{AT(\Omega)})=h_{0}([\widetilde{\mu}]_{T(\Omega)})=k_{0}([\widetilde{\mu}]_{T(\Omega)})=\|\widetilde{\mu}\|_{\infty}$. It is easy to see that $H_{0}([\mu]_{AT(\Omega)})=h_{0}([\widetilde{\mu}]_{T(\Omega)})=k_{0}([\widetilde{\mu}]_{T(\Omega)})=h^{*}(\widetilde{\mu})=\|\widetilde{\mu}\|_{\infty}$, if $\widetilde{\mu}$ is non-Strebel extremal in $\mu]_{AT({\Omega})}$. From [55, Lemma 4.1], we know that
@@ 5.2. Proof of Theorem 4
Hence we see from [8] or [4], [56] that $\mathbf{F}$ is bounded univalent in $\Delta$ and admits a quasiconformal extension to $\widehat{\mathbb{C}}$ when $\varepsilon$ small enough. Now, we let $\varepsilon$ small enough so that $\mathbf{F}$ belongs to $\mathcal{S}_{q}$. On the other hand, we note that, for any $r\in(0,1)$,
@@
For any $\tau\in\mathbb{C}$, note that $N_{{\bf F}}=(1+\varepsilon)N_{f}$ and $\beta_{\mathbf{F}}(\tau)$$\beta_{f}(\tau+\varepsilon\tau)$ in the above proof of Theorem 4, we see that, for any $f\in\mathcal{S}_{q}$, there is always a $\phi\in\partial\mathcal{T}$ such that $\phi=aN_{f},a>1$ and $\beta_{f_{\phi}}(\tau)\geq\beta_{f}(\tau)$. This means that
@@ 6. Integral means spectra of the univalent rational functions
In this section, we study the integral means spectra of univalent rational functions in $\Delta$. In the rest of the paper, the notation $P_{n}(z)$ denotes a polynomial of degree at most $n$, $n\in\mathbb{N}\cup\{0\}$. In particular, $P_{n}$ is a complex constant when $n=0$. We denote by $\mathcal{R}$ the class of all rational functions $R$ in $\Delta$ with $R(0)=R^{\prime}(0)-1=0$ and which have no poles in $\Delta$. Let $\Omega$ be a subset of the complex plane, we say $R\in\mathcal{R}$ has no zeros, or one zero in $\Omega$ if the equation $R(z)=0$ has no roots, or one root in $\Omega$, respectively. We will say $R\in\mathcal{R}$ has no critical points, or one critical point if the equation $R^{\prime}(z)=0$ has no roots, or one root in $\Omega$, respectively. A rational function $R\in\mathcal{R}$ can be uniquely represented as
@@
6.1. Univalent rational functions and the main result of this section
@@
As we shall see later, the integral means spectra of functions $R\in\mathcal{U}_{R}$ that have no critical points in $\mathbb{T}$, are different from those of functions $R$ that have at least one critical point in $\mathbb{T}$. We will list some examples of univalent rational functions $R(z)$ which has no critical points, or has at least one critical point in $\mathbb{T}$. We first consider the univalent polynomials.
@@
Take $P_{1}(z)=z$, then $P^{\prime}_{1}(z)=1$ so that $P_{1}$ has no critical points in $\mathbb{T}$.
@@
Take $P_{2}(z):=z-\frac{1}{2}z^{2}$, then $P^{\prime}_{2}(z)=1-z$ so that $P_{2}$ has one critical point in $\mathbb{T}$.
@@
Take $P_{3}(z):=z-\frac{1}{3}z^{3}$, then $P^{\prime}_{3}(z)=(1-z)(1+z)$ so that $P_{3}$ has two critical points in $\mathbb{T}$.
@@
Take $P_{5}(z)=z-\frac{1}{6}z^{3}-\frac{1}{10}z^{5}$, then $P^{\prime}_{5}(z)=-\frac{1}{2}(z^{2}-1)(z^{2}+2)$ so that $P_{5}$ has two critical points in $\mathbb{T}$ and two other ones not in $\mathbb{T}$.
@@
$P_{1},P_{2},P_{3},P_{4}$ are all contained in $\mathcal{U}_{P}\cap\mathcal{L}_{I}$.
@@
Take $R_{1}(z)=\frac{z}{1-z},$ then $R_{1}^{\prime}(z)=\frac{1}{(1-z)^{2}}$ so that $R_{1}$ has no critical points in $\mathbb{T}$.
@@
Take $R_{2}(z)=\frac{z^{2}+3z}{3(1-z)},$ then $R_{2}^{\prime}(z)=-\frac{(z-3)(z+1)}{3(1-z)^{2}}$ so that $R_{2}$ has one critical point in $\mathbb{T}$.
@@
Take $R_{3}(z)=\frac{1}{(1-z)^{2}},$ then $R_{3}^{\prime}(z)=\frac{2}{(1-z)^{3}}$ so that $R_{3}$ has no critical points in $\mathbb{T}$.
@@
Take $R_{4}(z)=\kappa(z):=\frac{z}{(1-z)^{2}},$ here $\kappa$ is the famous Keobe function, then $R_{4}^{\prime}(z)=\frac{1+z}{(1-z)^{3}}$ so that $R_{4}$ has one critical point in $\mathbb{T}$.
@@
When $R\in\mathcal{U}_{R}$. If $R$ has a critical point $z_{1}$ in $\mathbb{T}$, because $R$ is univalent, it follows that
@@
Generally, if $R$ has at least one critical point in $\mathbb{T}$, then $R^{\prime}(z)$ can be written as the following unique form.
@@
Here, $s\in\mathbb{N}$, and $z_{j},j=1,2,\cdots,s$ are all distinct critical points of $R$ in $\mathbb{T}$. The numerator and denominator of $R^{\prime}$ have no common factors except $1$. $P_{\bf n}(z)$ has no zeros in $\mathbb{T}$, and $P_{\bf m}(z)$ has no zeros in $\Delta$.
@@
(6.4) $R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}:=\Pi(z)\frac{P_{\bf n}(z)}{P_{\bf m}(z)}.$
Here $s\in\mathbb{N}\cup\{0\}$. When $s\geq 1$, $z_{j},j=1,2,\cdots,s$ are all distinct critical points in $\mathbb{T}$ of $R$. The numerator and denominator of $R^{\prime}$ have no common factors except $1$. $P_{\bf n}(z)$ has no zeros in $\mathbb{T}$, and $P_{\bf m}(z)$ has no zeros in $\Delta$. The case $s=0$ means that $\Pi(z)\equiv 1$ and $R$ has no critical points in $\mathbb{T}$.
@@
Then from
$R^{\prime}(z)=\frac{P^{\prime}_{n}(z)P_{m}(z)-P_{n}(z)P^{\prime}_{m}(z)}{[P_{m}(z)]^{2}}=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)},$
we see that $P_{\bf m}(z)$ has no zeros in $\overline{\Delta}.$ This means that there exists a neighborhood of $\mathbb{T}$, i.e., an annulus centered at the origin containing $\mathbb{T}$, in which $R^{\prime}$ has no poles and $z_{1},z_{2},\dots,z_{s}$ are all zeros of $R^{\prime}$. In particular, $R^{\prime}$ has no zeros in this neighborhood when $s=0$.
@@
We will show that the Brennan conjecture B(2)=1B(-2)=1 is true for the class $\mathcal{U}_{R}$. Actually, we shall completely determine the generalized integral means spectra of functions belonging to the class $\mathcal{R}_{O}\cap({\mathcal{L}_{I}}\cup{\mathcal{L}_{II}}\cup{\mathcal{L}_{III}})$ for all real $\tau$. Let $\tau\in\mathbb{R}$. For each $R\in\mathcal{R}$, since $R$ has only finitely many critical points, we know that there is a constant $r_{R}\in(0,1)$ such that $R^{\prime}(z)\neq 0$ for all $z\in\{z:r_{R}<|z|<1\}$. We define the generalized integral means spectrum ${\widetilde{\beta}}_{R}$ of $R$ as
(6.7) ${\widetilde{\beta}}_{R}(\tau):=\lim\limits_{r\rightarrow 1^{-}}\sup\limits_{r\in(r_{R},1)}\frac{\log\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta}{|\log(1-r)|}.$
@@
Note that this definition is the same as in (1.1) when $R$ belongs to $\mathcal{U}_{R}$.
@@
$({\bf b_{1}})$ if $R$ has no critical points in $\mathbb{T}$, then ${\widetilde{\beta}}_{R}(\tau)=0$,
$({\bf b_{2}})$ if $R$ has at least one critical point in $\mathbb{T}$, then
@@
Let $\varkappa>0$, $r\in(0,1)$, $n\in\mathbb{N}$ and $z_{j}=e^{i\arg z_{j}}$, $j=1,2,\cdots,n$ be distinct points in $\mathbb{T}$. We assume that $g$ is an analytic function in $\Delta$ and satisfies that,
@@
$({\bf b_{1}})$ If $R$ has at east one critical point in $\mathbb{T}$, we let, as in (6.4),
$R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}:=\prod_{j=1}^{s}(z-z_{j})g(z).$
From Remark 6.16, for $R\in\mathcal{R}_{O}\cap\mathcal{L}_{I}$, or $R\in\mathcal{R}_{O}\cap\mathcal{L}_{II}$, or $R\in\mathcal{R}_{O}\cap\mathcal{L}_{III}$, we see that
@@
$({\bf b_{2}})$ If $R$ has no critical points in $\mathbb{T}$. Then, it is easy to see from the definition (6.7) of the generalized integral means spectrum that ${\widetilde{\beta}}_{R}(\tau)=0$ for any $\tau<0$.
@@
Here, the numerator and denominator of $Q_{1}(z)$ have no common factors except $1$. Then, from part (b) of Remark 6.16, we can find constants $M>0$, $r_{0}\in(r_{R},1)$, $\eta_{j}\in(0,\frac{\pi}{2l})$, $M_{j,1}>0,M_{j,2}>0$, $j=1,2,\cdots,l$, such that

### v10

@@ Abstract.
In this paper we investigate the complex exponential integral means spectra of univalent functions in the unit disk. We show that all integral means spectrum (IMS) functionals for complex exponents on the universal Teichmüller space, the closure of the universal Teichmüller curve, and the universal asymptotic Teichmüller space are continuous. We also show that the complex exponential integral means spectrum of any univalent function admitting a quasiconformal extension to the extended complex plane is strictly less than the universal integral means spectrum. These extend some related results in our recent work [38]. Here we employ a different and more direct approach to prove the continuity of the IMS functional on the universal asymptotic Teichmüller space. Additionally, we completely determine the integral means spectra of all univalent rational functions in the unit disk. As a consequence, we show that the Brennan conjecture is true for this class of univalent functions. Finally, we present some remarks and raise some problems and conjectures regarding IMS functionals on Teichmüller spaces, univalent rational functions, and a multiplication operator whose norm is closely related to the Brennan conjecture.
@@
Complex exponential integral means spectrum of univalent function, universal integral means spectrum, Brennan conjecture, univalent function with a quasiconformal extension, univalent rational function, univalent polynomial, universal Teichmüller space, universal asymptotic Teichmüller space, universal Teichmüller curve
@@ 1. Introduction
For real $\tau$, the integral means measure the boundary expansion and compression associated with a given conformal map. Allowing $\tau$ to be a complex exponent, we should also take into account the rotation.
The universal integral means spectra $B(\tau)$ and $B_{b}(\tau)$ are defined as
@@
Because of its important relation to the fine properties of harmonic measures in the complex plane, see [44], the universal integral means spectra of univalent functions have been studied extensively in recent years. It is an important and difficult problem to find out the exact values of the universal integral means spectra $B(\tau)$ and $B_{b}(\tau)$. We shall first review some main known results and open problems for real $\tau$ in this topic. It was first observed by Makarov in [44] that
@@
(1) For fixed $\tau\in\mathbb{C}$ with $\tau\neq 0$, what topology can be endowed upon the class $\mathcal{S}_{q}$ such that the integral means spectrum $\beta_{f}(\tau)$ is continuous on $\mathcal{S}_{q}$? (2) For given $\tau\neq 0$, does there exist at least one extremal function for $B_{b}(\tau)$? (3) If the extremal functions for $B_{b}(\tau)$ exist, in which subset of $\mathcal{S}$ do they lie?
Here and later, we say a function $f\in\mathcal{S}$ is an extremal function for $B_{b}(\tau)$ if $\beta_{f}(\tau)=B_{b}(\tau)$. In the recent work [38], we have studied Problem 1.11 for real $\tau$. It has been pointed out that the functional $I:f\rightarrow\beta_{f}(\tau),f\in\mathcal{S}_{q}$, is not continuous on $\mathcal{S}_{q}$ under the locally uniformly convergence topology, see [38, Remark 5.8]. In [38], we introduced and studied the integral means spectrum (IMS) functionals on the Teichmüller spaces. It has been proved in [38] that all IMS functionals on the universal Teichmüller space, the closure of the universal Teichmüller curve, and the universal asymptotic Teichmüller space are continuous. It was also shown in [38] that the real exponential integral means spectrum of any univalent function admitting a quasiconformal extension to $\widehat{\mathbb{C}}$ is strictly less than the universal integral means spectrum. This result means that, for $\tau\in\mathbb{R}$ with $\tau\neq 0$, the extremal functions for $B_{b}(\tau)$ cannot be from the class $\mathcal{S}_{q}$. The first main purpose of this paper is to extend these results to the complex setting. We shall prove that
@@ 4. Proof of Theorem 1 and 2
Note that $d_{AT}([\mu]_{AT},[\nu]_{AT})\leq d_{T}([\mu]_{T},[\nu]_{T})$ for any $\mu,\nu\in M(\Delta^{*})$, we see that the statement that $I_{AT}$ is continuous on $AT$ implies Theorem 1. Hence we only need to prove Theorem 2. By studying the Pre-Schwarzian derivative model of the universal asymptotic Teichmüller space, we have proved in [38] that Theorem 2 holds for real $\tau$. We will use a more direct way to show that Theorem 2 holds for all complex numbers $\tau$. More precisely, we will prove Theorem 2 by using some arguments from the theory of Teichmüller spaces and some results from the extremal theory of quasiconformal mappings.
@@ 4.1. Isomorphism mapping between Teichmüller spaces
Here $\chi=\frac{\overline{\partial f_{\omega}}}{\partial f_{\omega}}$. We will next write $\mu^{\sharp}(\zeta)=\sharp(\mu)(\zeta)$. Then, from $f_{\mu}=f_{\mu^{\sharp}}\circ f_{\omega}$, we obtain
@@ 4.2. Extremal theory of quasiconformal mappings
It is known that $H_{0}([\mu]_{AT(\Omega)})=h_{0}([\mu]_{T(\Omega)})$. We say some $\widetilde{\mu}\in[\mu]_{AT({\Omega})}$ is non-Strebel extremal in $\mu]_{AT({\Omega})}$ if it satisfies that $H_{0}([\mu]_{AT(\Omega)})=h_{0}([\widetilde{\mu}]_{T(\Omega)})=k_{0}([\widetilde{\mu}]_{T(\Omega)})=\|\widetilde{\mu}\|_{\infty}$. It is easy to see that $H_{0}([\mu]_{AT(\Omega)})=h_{0}([\widetilde{\mu}]_{T(\Omega)})=k_{0}([\widetilde{\mu}]_{T(\Omega)})=h^{*}(\widetilde{\mu})=\|\widetilde{\mu}\|_{\infty}$, if $\widetilde{\mu}$ is non-Strebel extremal in $\mu]_{AT({\Omega})}$. From [56, Lemma 4.1], we know that
@@ 5.2. Proof of Theorem 4
Hence we see from [8] or [4], [57] that $\mathbf{F}$ is bounded univalent in $\Delta$ and admits a quasiconformal extension to $\widehat{\mathbb{C}}$ when $\varepsilon$ small enough. Now, we let $\varepsilon$ small enough so that $\mathbf{F}$ belongs to $\mathcal{S}_{q}$. On the other hand, we note that, for any $r\in(0,1),\theta\in(-\pi,\pi]$,
@@
Note that for any $f\in\mathcal{S}_{q}$, there is always an $F\in\mathcal{S}$ such that $N_{F}=\phi\in\partial\mathcal{T}$ and $\phi=aN_{f},a>1$. We let $\mathbf{h}=F\circ f^{-1}$. Then $\mathbf{h}$ is univalent in $f(\Delta)$ and
$(a-1)N_{f}=N_{F}(z)-N_{f}(z)=\frac{\mathbf{h}^{\prime\prime}\circ f(z)}{\mathbf{h}^{\prime}\circ f(z)}\cdot f^{\prime}(z).$
Integrate both sides of the above equation from 0 to $z$, we obtain that
$(a-1)\log f^{\prime}(z)=\log\mathbf{h}^{\prime}\circ f(z).$
Hence $\mathbf{h}^{\prime}\circ f(z))=[f^{\prime}(z)]^{a-1}$. It follows that for any $r\in(0,1),\theta\in(-\pi,\pi]$,
$[F^{\prime}(re^{i\theta})]^{\tau}$ $=$ $[\mathbf{h}^{\prime}\circ f(re^{i\theta})]^{\tau}\cdot[f^{\prime}(re^{i\theta})]^{\tau}=[f^{\prime}(re^{i\theta})]^{a\tau},\tau\in\mathbb{C}.$
From Remark 5.2, we have $\beta_{F}(\tau)=\beta_{f}(a\tau)\geq\beta_{f}(\tau)$. This implies that
@@ 6. Integral means spectra of the univalent rational functions
In this section, we study the integral means spectra of univalent rational functions in the unit disk $\Delta$. Since directly studying the integral means spectrum of univalent rational functions in $\Delta$ is inconvenient, we enlarge the scope to the general class of rational functions in $\Delta$. The rational functions, having only finitely many critical points in $\Delta$, allows for the definition of a generalized integral means spectrum, see (6.8) below. In particular, this generalized spectrum coincides with the classical one for the univalent rational functions. We then completely determine the generalized integral means spectra for a certain class of rational functions, thereby proving that univalent rational functions in the unit disk satisfy the Brennan conjecture.
In the rest of the paper, the notation $P_{n}(z)$ denotes a polynomial of degree at most $n$, $n\in\mathbb{N}\cup\{0\}$. In particular, $P_{n}$ is a complex constant when $n=0$. We denote by $\mathcal{R}$ the class of all rational functions $R$ in $\Delta$ with $R(0)=R^{\prime}(0)-1=0$ and which have no poles in $\Delta$. Let $\Omega$ be a subset of the complex plane, we say $R\in\mathcal{R}$ has no zeros, or one zero in $\Omega$ if the equation $R(z)=0$ has no roots, or one root in $\Omega$, respectively. We will say $R\in\mathcal{R}$ has no critical points, or one critical point if the equation $R^{\prime}(z)=0$ has no roots, or one root in $\Omega$, respectively. A rational function $R\in\mathcal{R}$ can be uniquely represented as
@@
6.1. Certain subclasses of rational functions and the main result of this section
@@
As we shall see later, the integral means spectra of functions $R\in\mathcal{U}_{R}$ that have no critical points on $\mathbb{T}$, are different from those of functions $R$ that have at least one critical point on $\mathbb{T}$. We will list some examples of univalent rational functions $R(z)$ which has no critical points, or has at least one critical point on $\mathbb{T}$. We first consider the univalent polynomials.
@@
Take $P_{1}(z)=z$, then $P^{\prime}_{1}(z)=1$ so that $P_{1}$ has no critical points on $\mathbb{T}$.
@@
Take $P_{2}(z):=z-\frac{1}{2}z^{2}$, then $P^{\prime}_{2}(z)=1-z$ so that $P_{2}$ has one critical point on $\mathbb{T}$.
@@
Take $P_{3}(z):=z-\frac{1}{3}z^{3}$, then $P^{\prime}_{3}(z)=(1-z)(1+z)$ so that $P_{3}$ has two critical points on $\mathbb{T}$.
@@
Take $P_{5}(z)=z-\frac{1}{6}z^{3}-\frac{1}{10}z^{5}$, then $P^{\prime}_{5}(z)=-\frac{1}{2}(z^{2}-1)(z^{2}+2)$ so that $P_{5}$ has two critical points on $\mathbb{T}$ and two other ones not on $\mathbb{T}$.
@@
$P_{1},P_{2},P_{3},P_{5}$ are all contained in $\mathcal{U}_{P}\cap\mathcal{L}_{I}$.
@@
Take $R_{1}(z)=\frac{z}{1-z},$ then $R_{1}^{\prime}(z)=\frac{1}{(1-z)^{2}}$ so that $R_{1}$ has no critical points on $\mathbb{T}$.
@@
Take $R_{2}(z)=\frac{z^{2}+3z}{3(1-z)},$ then $R_{2}^{\prime}(z)=-\frac{(z-3)(z+1)}{3(1-z)^{2}}$ so that $R_{2}$ has one critical point on $\mathbb{T}$.
@@
Take $R_{3}(z)=\frac{1}{(1-z)^{2}},$ then $R_{3}^{\prime}(z)=\frac{2}{(1-z)^{3}}$ so that $R_{3}$ has no critical points on $\mathbb{T}$.
@@
Take $R_{4}(z)=\kappa(z):=\frac{z}{(1-z)^{2}},$ here $\kappa$ is the famous Koebe function, then $R_{4}^{\prime}(z)=\frac{1+z}{(1-z)^{3}}$ so that $R_{4}$ has one critical point on $\mathbb{T}$.
@@
When $R\in\mathcal{U}_{R}$. If $R$ has a critical point $z_{1}$ on $\mathbb{T}$, because $R$ is univalent, it follows that
@@
Generally, if $R$ has at least one critical point on $\mathbb{T}$, then $R^{\prime}(z)$ can be written as the following unique form.
@@
Here, $s\in\mathbb{N}$, and $z_{j},j=1,2,\cdots,s$ are all distinct critical points of $R$ on $\mathbb{T}$. The numerator and denominator of $R^{\prime}$ have no common factors except $1$. $P_{\bf n}(z)$ has no zeros on $\mathbb{T}$, and $P_{\bf m}(z)$ has no zeros in $\Delta$.
@@
(6.4) $R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}:=\frac{\Pi(z)P_{\bf n}(z)}{P_{\bf m}(z)}.$
Here $s\in\mathbb{N}\cup\{0\}$. When $s\geq 1$, $z_{j},j=1,2,\cdots,s$ are all distinct critical points on $\mathbb{T}$ of $R$. The numerator and denominator of $R^{\prime}$ have no common factors except $1$. $P_{\bf n}(z)$ has no zeros on $\mathbb{T}$, and $P_{\bf m}(z)$ has no zeros in $\Delta$. The case $s=0$ means that $\Pi(z)\equiv 1$ and $R$ has no critical points on $\mathbb{T}$. Note that we allow $P_{\bf n}(z)$ to have zeros in the unit disk.
@@
Then
(6.5) $R^{\prime}(z)=\frac{P^{\prime}_{n}(z)P_{m}(z)-P_{n}(z)P^{\prime}_{m}(z)}{[P_{m}(z)]^{2}}=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}$
Note that the numerator and denominator of the last rational function in (6.5) have no common factors except $1$ so that $P_{\bf m}(z)$ has no zeros in $\overline{\Delta}.$ This means that there exists a neighborhood of $\mathbb{T}$, i.e., an open annulus centered at the origin containing $\mathbb{T}$, in which $R^{\prime}$ has no poles and $z_{1},z_{2},\dots,z_{s}$ are all zeros of $R^{\prime}$. In particular, $R^{\prime}$ has no zeros in this neighborhood when $s=0$.
@@
Next, we will mainly focus on the rational functions whose derivatives have the forms (6.5) or (6.16) or (6.16). These forms shall play an important role in our proof of the main result of this section below.
We will show that the Brennan conjecture B(2)=1B(-2)=1 is true for the class $\mathcal{U}_{R}$. Actually, we shall completely determine the generalized integral means spectra of functions belonging to the class $\mathcal{R}_{O}\cap({\mathcal{L}_{I}}\cup{\mathcal{L}_{II}}\cup{\mathcal{L}_{III}})$ for all real $\tau$. Let $\tau\in\mathbb{R}$ and $f$ be an analytic function in $\Delta$. If there is a constant $r_{f}\in(0,1)$ such that $f^{\prime}(z)\neq 0$ for all $z\in\{z:r_{f}<|z|<1\}$, then we define the generalized integral means spectrum ${\widetilde{\beta}}_{f}$ of $f$ as
(6.8) ${\widetilde{\beta}}_{f}(\tau):=\limsup\limits_{r\rightarrow 1^{-}}\frac{\log\int_{-\pi}^{\pi}|f^{\prime}(re^{i\theta})|^{\tau}d\theta}{|\log(1-r)|}.$
Since $f^{\prime}$ has no zeros for $r>r_{f}$, the quantity $|f^{\prime}(re^{i\theta})|^{\tau}$ is well-defined for all $r$ sufficiently close to $1$, and thus the $\limsup$ is meaningful. For each $R\in\mathcal{R}$, since $R$ has only finitely many critical points, we know that there is a constant $r_{R}\in(0,1)$ such that $R^{\prime}(z)\neq 0$ for all $z\in\{z:r_{R}<|z|<1\}$. Hence we can define the generalized integral means spectrum ${\widetilde{\beta}}_{R}$ of $R$ as in (6.8).
@@
Note that, for any $\tau\in\mathbb{R}$, this definition is the same as in (1.1) when $R$ belongs to $\mathcal{U}_{R}$.
@@
$({\bf b_{1}})$ if $R$ has no critical points on $\mathbb{T}$, then ${\widetilde{\beta}}_{R}(\tau)=0$,
$({\bf b_{2}})$ if $R$ has at least one critical point on $\mathbb{T}$, then
@@
Let $\varkappa>0$, $r\in(0,1)$, $n\in\mathbb{N}$ and $z_{j}=e^{i\arg z_{j}}$, $j=1,2,\cdots,n$ be distinct points on $\mathbb{T}$. We assume that $g$ is an analytic function in $\Delta$ and satisfies that,
@@
$({\bf b_{1}})$ If $R$ has at east one critical point on $\mathbb{T}$, we let, as in (6.4),
$R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}:=g(z)\prod_{j=1}^{s}(z-z_{j}).$
In view of (6.5), (6.16) and (6.16) in the Remark 6.16, we see that for $R\in\mathcal{R}_{O}\cap\mathcal{L}_{I}$, or $R\in\mathcal{R}_{O}\cap\mathcal{L}_{II}$, or $R\in\mathcal{R}_{O}\cap\mathcal{L}_{III}$, we have
@@
$({\bf b_{2}})$ If $R$ has no critical points on $\mathbb{T}$. Then, we see that $R$ is bounded above and below in a neighborhood of $\mathbb{T}$. Hence we obtain from the definition (6.8) of the generalized integral means spectrum that ${\widetilde{\beta}}_{R}(\tau)=0$ for any $\tau<0$.
@@
Here, the numerator and denominator of the rational function $Q_{1}(z)$ have no common factors except $1$.
Then, from part (b) of Remark 6.16, we can find constants $M>0$, $r_{0}\in(r_{R},1)$, $\eta_{j}\in(0,\frac{\pi}{2l})$, $M_{j,1}>0,M_{j,2}>0$, $j=1,2,\cdots,l$, such that